Proof whether the following operations are inner product operations:

⟨x, y⟩ = 2x1y1 − x1y2 − x2y1 + 2x2y2, x=(x1, x2), y=(y1, y2)



1
Expert's answer
2021-11-22T03:42:38-0500

 "\u27e8x, y\u27e9" satisfies the following properties:


Linearity:

"(ax+by,z)=a(x,z)+b(y,z)"

"(ax+by,z)= 2(ax_1+by_1)z_1 \u2212 (ax_1+by_1)z_2 \u2212 (ax_2+by_2)z_1 + 2(ax_2+by_2)z_2="

"=a( 2x_1z_1 \u2212 x_1z_2 \u2212 x_2z_1 + 2x_2z_2)+b( 2y_1z_1 \u2212 y_1z_2 \u2212 y_2z_1 + 2y_2z_2)"

"a(x,z)+b(y,z)=a( 2x_1z_1 \u2212 x_1z_2 \u2212 x_2z_1 + 2x_2z_2)+b( 2y_1z_1 \u2212 y_1z_2 \u2212 y_2z_1 + 2y_2z_2)"


Symmetric Property:

"(x,y)=(y,x)"


Positive Definite Property:

"⟨x, x⟩ = 2x_1x_1 − x_1x_2 − x_2x_1 + 2x_2x_2=(x_1-x_2)^2+x_1^2+x_2^2\ge0"


so,

"\u27e8x, y\u27e9" is inner product